Two jets leave Dallas and fly in opposite directions. One is flying 50 mph faster than the other. After 2 hours they are 2500 miles apart. Find the speed of each jet.
Answer by orca(409)

Let x be the speed of the slower jet.
Then the speed of the faster one is x + 50.
The distance that separating them at the end of 2 hours can be expressed in terms of x as:
2x + 2(x + 50)
As the separating distance is 2500, we have
2x + 2(x + 50) = 2500
Solving for x, we have
2x + 2x +100 = 2500
4x = 2400
x = 600
So the speed of the slower jet is 600 miles per hour
the speed of the fast jet is x + 50 = 650 miles per hour.